Meeting Notes:

1. General status updates/Schedule review.
Eric -- artdaq v3_06_01 is almost ready for release, then otsdaq release will be next, then Mu2e test stand. Also worked a bit with Rick pcie_linux_kernel_module.
John -- All 88 production timing cards have been tagged and entered in spreadsheet, along with 5 STM-only v2 timing cards. 50 cassettes have tags with stickers applied. Next will populate Fireflies onto timing cards - Fireflies have been scanned into spreadsheet and will record the pairing. FPGAs are in, will be paired with timing cards.
Andrei -- here to listen
Micol -- worked with Monica on DRAC and was able to see data replies from Tracker board. Now working on Loopback, but stuck with Libero being difficult.
Antonio -- now able to read live information from EPICS and archived values from EPICS. Working on optimization of ots as server and web-GUI as client.
Rick -- Working hardware event building - making good progress. Switching his office test stand to new production timing cards, including CFO. Finding issue with CFO - could remove clock chip or use old boards. CRV Lei says working on a firmware timing issue.
Vivian -- Helped get data off tracker board with Micol and Monica - sees issues in data. More testing needed. Monica working on new firmware version.
Ron --
Sophie -- Heard from Erdem, not much progress. Stefan and Oliver are here right now - Erdem should pull latest ROC changes.
Ryan -- We have all our productions hardware (except detector optical links).
Glenn -- Mu2e CRV keeping KSU busy - test stand going pretty well. Sten is changing frequency of onboard oscillator. We should use standard EPICS CSS Archiver in postgres (archiver appliance from SLAC is too raw). Currently running on mu2edaq12.
Craig -- Sten has been wrestling with timing and clock distribution for 3-4 weeks.
Nam -- Resumed working on mongo db usage.. Seeing issues with table loading.
Bertrand -- How will we backup EPICS? Glenn is going to ask SCD to setup “read-only” offline replication for EPICS db.

2. Status of DTC to CRV ROC interface.
3. Status of EPICS on DAQ cluster.
4. Status or core ROC firmware testing.